Please help me solve these Trigonometric Addition Formulas:
cos2xcos3x+sin2xsin3x
sin5xcosx-cos5xsinx

Answer by lwsshak3(11628)

cos2xcos3x+sin2xsin3x
Identity: cos(s-t)=cos s cos t+sins sin t
cos2xcos3x+sin2xsin3x
=cos(2x-3x)
=cos(-x)
=cosx (cos is an even function)
..
sin5xcosx-cos5xsinx
Identity: sin(s-t)=sin s cos t-cos s sin t
sin5xcosx-cos5xsinx
=sin(5x-x)
=sin (4x)